### 12 Dec 2023 v2

@@ 2.3 Demonstration of equivalence
We note from this that $X_{2}$ is a precision variable (related strongly to the outcome but weakly to the probability of inclusion in the sample), and $X_{4}$ is a bias variable (related strongly to both). According to Little & Vartivarian (2005), models with $X_{4}$ and not $X_{2}$ should be strongly preferred over models with $X_{2}$ and not $X_{4}$.
As we are focusing on a binary outcome, we create observed binary outcomes $y_{i}$ for every individual in the population using $\mbox{Pr}(y_{i}=1)$. We also discretize the $X^{(k}$ variables into $5$ groups of equal range, which we use to model.
We simulate with a population size of $20,000$, twice the size of Kuh et al. (2022). We sample from the population firstly so that every cell has one observation and the secondly with probability $\pi_{i}$. The sample size is held constant at $1,000$. The population creation and sampling is repeated $100$ times. As the outcome is binary, we choose to fit a binomial likelihood of counts per cell $j$. We also only fit a subset of the models used by Kuh et al. (2022), focusing on the following models:
  • Full model: $\mbox{Pr}(y_{j}=1|n_{j})=\operatorname{logit}^{-1}(\beta_{0}+\alpha^{(X_{1})}_{j}+\alpha^{(X_{2})}_{j}+\alpha^{(X_{3})}_{j}+\alpha^{(X_{4})}_{j}),$
  • Precision variable model: $\mbox{Pr}(y_{j}=1|n_{j})=\operatorname{logit}^{-1}(\beta_{0}+\alpha^{(X_{1})}_{j}+\alpha^{(X_{2})}_{j}+\alpha^{(X_{3})}_{j}),$
  • Bias variable model: $\mbox{Pr}(y_{j}=1|n_{j})=\operatorname{logit}^{-1}(\beta_{0}+\alpha^{(X_{1})}_{j}+\alpha^{(X_{3})}_{j}+\alpha^{(X_{4})}_{j}),$
  • Nuisance variable model: $\mbox{Pr}(y_{j}=1|n_{j})=\operatorname{logit}^{-1}(\beta_{0}+\alpha^{(X_{1})}_{j}+\alpha^{(X_{3})}_{j}),$
where $\alpha^{k}_{j}\sim\operatorname{normal}(0,\sigma^{(k)}$), $\sigma^{(k)}\sim t(3,0,2.5)$ and $\beta_{0}\sim t(3,0,2.5)$, which are the standard priors in the brms R package (Bürkner, 2017, 2018).
Figure 1: Comparison of a direct score of the MRP estimate (using known population truth, $y$-axis) against an indirect score of the MRP estimate (using the known population truth for each cell, $x$-axis). Colour of point represents the different models fitted, whilst different points within this represent different simulation iterations. The black line represents an identity line. As the points all fall either directly or with a small amount of variation, we find evidence that this decomposition of the scoring rule is valid.
@@
    Create a sample from this population so that each cell has at least one individual observed, and then according to $\pi_{i}$.
@@
      An estimate for each cell and compare to the true $\mbox{Pr}(Y_{j}=1)$ to calculate the cellwise CRPS and squared error as per (10) and (6).
Figure 1 plots the score for our population mean estimated with MRP against the score calculated through the cellwise error. Up to a small amount of noise, these two scores are equivalent. This provides confirmation for our proposal that the score for a population mean estimate can be decomposed into cell estimates.
3 Estimating MRP goodness using sample goodness
3.1 Sample as proxy
Continuing with the previous example, we aim to score our model without population truth using the observed values in each cell. We first consider a scenario where at least one individual in every cell is observed. We continue this assumption in Section 4 where we consider subpopulation estimates. In Section 4 we relax this assumption to consider techniques when not all cells are observed in the sample.
The results from Section 2 are a useful proof of the equivalence of these two formulations, but it is not directly clear how they might be used in a real world setting where the truth is not known. One example is to implement an approximation of (6) and (10), so that instead of using the population truth we instead approximate the population truth using a cellwise sample estimate which we denote $\mbox{Pr}(y_{i\in j}=1|n_{j})$. In these adaptions, we replace the population truth with this value:
$\widehat{\operatorname{Error}}_{\mathrm{MRP}}^{2}=\Bigg{(}\frac{\sum_{j=1}^{J}N_{j}(\mbox{Pr}(Y_{i\in j}=1|y,\textbf{x},\hat{\theta})-\mbox{Pr}(y_{i\in j}=1|n_{j},y_{j}))}{\sum_{j=1}^{J}N_{j}}\Bigg{)}^{2},$ (14)
$\widehat{\operatorname{CRPS}}_{\mathrm{MRP}}(y,n,\hat{\theta},\mathbf{x})=\frac{1}{2}\frac{1}{B}\sum_{b}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}\Big{(}\mbox{Pr}(Y_{i\in j}|y,\hat{\theta}^{b},\textbf{x})-\mbox{Pr}(Y_{i\in j}|y,\hat{\theta}^{\prime b},\textbf{x})\Big{)}}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}-\\ \frac{1}{B}\sum_{b}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}\Big{(}\mbox{Pr}(Y_{i\in j}|y,\hat{\theta}^{b},\textbf{x})-\mbox{Pr}(y_{i\in j}=1|n_{j},y_{j})\Big{)}}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}.$ (15)
When there is only one observation per cell, this probability estimate is either 00 or $1$ accordingly. Figure 2 shows this new estimate on the $x$-axis and the true score on the $y$-axis. In both plots, we see two things relative to Figure 1. Our estimates of the model score are noisier but retain the basic ordering of models previously observed—namely, the separation between models that contain the bias variable as a predictor and models that contain only the precision variable as a predictor. This is important as the standard sum of individual error does not maintain that separation (e.g., see Kuh et al., 2022 for sum of ELPD, and supplementary materials for mean of CRPS and squared error scores). However, the error as estimated in the sample is a considerable underestimate relative to the truth. This could be explained by using the sample without correcting for overfitting.
Figure 2: Comparison of a direct score of the MRP estimate (using known population truth, $y$-axis) against an indirect score (using the known sample truth for each cell, $x$-axis). Colour of point represents the different models fitted, whilst different points within this represent different simulation iterations. The black line represents the identity. The CRPS scores estimated with the sample are overall higher than the true score (where lower indicates a worse model) and the squared error scores estimated with the sample are overall lower than the true score (where higher indicates a worse model). This indicates that using the sample as a proxy for population truth underestimates the model error.
@@ 3.2 Brute-force cross validation
To resolve the underestimation of the magnitude of scores seen in the previous section, we consider a cross validation approach. There is a wide and established literature on cross validation for estimating prediction error from a sample. There are three potential schemes that could be used for this problem:
  1. 1.
@@
    $K$-fold cross validation: In this approach we take split the sample into folds. Then we leave the $k^{th}$ fold out to fit the model and estimate error using that fold. In the context of MRP, we would need to adjust the error estimated in this fold to population. This would require that at least $k$ observations present in every poststratification matrix cell (one for each fold). While this might seem plausible, the size and complexity of survey data means that this is unlikely, even if MRP itself doesn’t require that every cell in the poststratification matrix is observed.
  2. 2.
    Leave-one-cell-out cross validation: In this implementation we leave out one cell of the poststratification matrix, predict the probability using a model fit with the remaining data, and then estimate model error using an estimate of $\mbox{Pr}(y_{j}=1|n_{j})$. In this scenario, only one observation is required per cell because it can be composed using our proposed formulae.
  3. 3.
    Leave-one-level-out (LOLO) cross validation: In this implementation we leave out one entire level of a adjustment variable ($x^{(k)}$), fit the model without it and predict for all observations in that level (this is also referred to as leave-one-group-out in the cross validation literature). However, it is not clear which variable to focus on when multiple variables are in the model. This approach could be useful approach for subpopulation estimation if the subpopulation represents different levels of a variable; see Section 4.
From this set of potential approaches, we feel that option 2 best aligns with the decomposition of total MRP score into cellwise components. This is not simply a special case of a $K$-fold because the individuals scores in each cell are not summed, but instead applied in adaptions to our previous formulae. Specifically we propose replacing our estimate $\mbox{Pr}(Y_{i\in j}=1|y,\textbf{x},\hat{\theta})$ with $\mbox{Pr}(Y_{i\in j}=1|y_{i\notin j},\textbf{x},\hat{\theta})$, the probability of $y=1$ in that cell given that it was not used when the model was fit:
@@
$\widehat{\operatorname{Error}}_{\mathrm{MRP,LOCO}}^{2}=\Bigg{(}\frac{\sum_{j=1}^{J}N_{j}(\mbox{Pr}(Y_{i\in j}=1|y_{i\notin j},\textbf{x},\hat{\theta})-\mbox{Pr}(y_{i\in j}=1|n_{j},y_{j}))}{\sum_{j=1}^{J}N_{j}}\Bigg{)}^{2},$ (16)
@@
$\widehat{\operatorname{CRPS}}_{\mathrm{MRP,LOCO}}(y,n,\hat{\theta},\mathbf{x})=\frac{1}{2}\frac{1}{B}\sum_{b}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}\Big{(}\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}^{b},\textbf{x})-\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}^{\prime b},\textbf{x})\Big{)}}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}-\\ \frac{1}{B}\sum_{b}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}\Big{(}\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}^{b},\textbf{x})-\mbox{Pr}(y_{i\in j}=1|n_{j},y_{j})\Big{)}}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}.$ (17)
@@
This brute-force implementation of leave-one-cell-out (LOCO) is not particularly efficient in terms of time/computational resources, but it does allow us to directly investigate whether the underestimation of error present found previously can be eliminated with this approach. Figure 3 shows that the error estimated with a brute-force LOCO approach is indeed slightly larger than that estimated using the full sample to model and estimate error. However, this is still a considerable underestimate for the true error, as shown by Figure 4.
Figure 3: Comparison of the brute-force leave-one-cell-out approximation (LOCO, $y$-axis) against the the indirect score of the MRP estimate (using the known sample truth for each cell, $x$-axis). Colour of point represents the different models fitted, whilst different points within this represent different simulation iterations. The black line represents an identity line. The CRPS scores estimated brute force LOCO are overall lower than the true score (where lower indicates a worse model) and the squared error scores estimated with brute force LOCO are overall higher than the true score (where higher indicates a worse model). This indicates that using brute force LOCO at least partially corrects the score underestimate when using the sample sample as a proxy for population truth.
@@
Figure 4: Comparison of a direct score of the MRP estimate (using known population truth, $y$-axis) against an indirect score of the MRP estimate (using a brute-force leave-one-cell-out approach, $x$-axis). Colour of point represents the different models fitted, whilst different points within this represent different simulation iterations. The black line represents an identity line. The CRPS scores estimated with LOCO are overall higher than the true score (where lower indicates a worse model) and the squared error scores estimated with LOCO are overall lower than the true score (where higher indicates a worse model). This indicates that using LOCO as a proxy for population truth still underestimates the model error.
@@
Despite the underestimate of the magnitude of the error, we believe that this approach is still useful to practitioners who wish to evaluate the relative efficacy of different models when estimating the population mean. Unlike previous attempts, ordering is maintained. We advocate that the magnitude of the error be interpreted with caution.
@@ 3.3 Approximate cross validation
The magnitude of the error is not the only challenge for a brute-force approach. A practical issue is the sheer amount of time and/or computational power used to run a brute-force leave-one-cell-out approach. For example, in the first iteration of the previously described simulation study, the posterior inference had to be rerun $294$ times (once for each cell plus one for full data), and this example is in many ways a relatively small toy problem in terms of the size of the poststratification matrix. Many MRP examples involve poststratifying to an area level variable like postcode, which might have $100$ levels (Machalek et al., 2022). In this section, we apply an approximate LOO method based on Pareto-smoothed importance-sampling (PSIS; Vehtari et al., 2017, 2022).
PSIS-LOO (Vehtari et al., 2017, Eqn 10) estimates the sum expected log individual predictive density as
@@
where $w_{i}^{b}$ is the Pareto-smoothed importance weight for the $b^{th}$ posterior draw (denoted $s$ by Vehtari et al., 2017) and the $i^{th}$ individual. We modify this for our purposes by using other scoring metrics. We begin with substituting the error in each cell as before, with for shorthand we denote $\epsilon_{j}^{b}=\mbox{Pr}(Y_{i\in j}=1|y_{i\notin j},\textbf{x},\hat{\theta}^{b})-\mbox{Pr}(Y_{i\in j}=1|n_{j},y_{j})$. For each cell $j$ we can compute Pareto-smoothed weights $w_{j}^{b}$ for each posterior sample $b$ to estimate $\epsilon_{j}^{b}$ as outlined by Vehtari et al. (2017). Using these, we can approximate $\epsilon_{j}^{b}$ and substitute into (16) as
$\widehat{\operatorname{Error}}_{\mathrm{PSIS-LOCO,MRP}}^{2}=\Bigg{(}\frac{1}{N}\sum_{j=1}^{J}N_{j}\frac{\sum_{b=1}^{B}w_{j}^{b}\epsilon_{j}^{b}}{\sum_{b=1}^{B}w_{j}^{b}}\Bigg{)}^{2}.$ (18)
@@
It is slightly more complicated to achieve this with the CRPS, but we begin by denoting the $g(y_{i\notin j},\textbf{x},\hat{\theta}^{b},\hat{\theta}^{\prime b})=\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}^{b},\textbf{x})-\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}^{\prime b},\textbf{x})$ and $h(y_{i\notin j},\textbf{x},y_{j},n_{j},\hat{\theta}^{b})=\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}^{b},\textbf{x})-\mbox{Pr}(y_{i\in j}=1|n_{j},y_{j})$. This means we can write (17) more concisely as
$\begin{split}\widehat{\operatorname{CRPS}}_{\mathrm{MRP}}(y_{i\notin j},y_{j},\textbf{x},\hat{\theta}^{b},\hat{\theta}^{\prime b})=\frac{1}{2}\frac{1}{B}\sum_{b=1}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}g(y_{i\notin j},\textbf{x},\hat{\theta}^{b},\hat{\theta}^{\prime b})}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}-\\ \frac{1}{B}\sum_{b=1}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}h(y_{i\notin j},\textbf{x},y_{j},n_{j},\hat{\theta}^{b})}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}.\end{split}$
While it is possible to simply take an importance-weighted sum for the squared error, CRPS doesn’t have an amenable form. To apply the PSIS weights, which are the $b^{th}$ weight for the $j^{th}$ cell, we do importance resampling of the posterior draws for each cell using a stratified resampling algorithm (Kitagawa, 1996) using the $B$ PSIS weights for each cell. This procedure is implemented in the posterior R package (Bürkner et al., 2023). This produces $B$ resampled posterior draws $\hat{\theta}_{RW}$, which we permute to calculate PSIS reweighted $\hat{\theta}_{RW}^{\prime}$, and thus $g(y_{i\notin j},\textbf{x},\hat{\theta}_{RW}^{b},\hat{\theta}_{RW}^{\prime b})$ and $h(y_{i\notin j},\textbf{x},y_{j},n_{j},\hat{\theta}_{RW}^{b})$, to get
$\begin{split}\widehat{\operatorname{CRPS}}_{\mathrm{PSIS-LOCO,MRP}}(y_{i\notin j},y_{j},\textbf{x},\hat{\theta}^{b},\hat{\theta}^{\prime b})=\frac{1}{2}\sum_{b=1}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}g(y_{i\notin j},\textbf{x},\hat{\theta}_{RW}^{b},\hat{\theta}_{RW}^{\prime b})}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}-\\ \sum_{b=1}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}h(y_{i\notin j},\textbf{x},y_{j},n_{j},\hat{\theta}_{RW}^{b})}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}.\end{split}$ (19)
Figure 5 shows that the approximate PSIS-LOCO scores are approximately equivalent to the brute-force LOCO scores.
Figure 5: Comparison of the fast Pareto-smoothed importance sampling approximation ($y$-axis) against the brute-force MRP LOCO estimate ($x$-axis). Colour of the point represents the different models fitted, whilst different points within this represent different simulation iterations. The black line represents an identity line. The strong relationship indicates that this approximation is highly efficient, with approximate methods closely following brute-force methods, with a small amount of noise.
@@
When we followed the traditional approach of comparing the mean of cell scores to the true mean of cell scores, the leave-one-cell-out approach appeared to slightly over estimate the magnitude of error (see supplementary), suggesting that the observed underestimation is a feature of our particular score decomposition.

### 12 Apr 2024 v3

@@ 3.1 Simulation Design
Here, $X_{2}$ is a precision variable (related strongly to the outcome but weakly to the probability of inclusion in the sample), and $X_{4}$ is a bias variable (related strongly to both). According to Little & Vartivarian (2005), models with $X_{4}$ and not $X_{2}$ should be strongly preferred over models with $X_{2}$ and not $X_{4}$.
We create binary observations $Y_{i}$ for every individual in the population using $\mbox{Pr}(Y_{i\in j}=1)$. We also discretize the $X_{k}$ variable into $5$ groups of equal range.
We simulate with a population size of $20,000$, twice the size of Kuh et al. (2023). We sample from the population by first ensuring every cell has one observation and then sampled the remaining obsevations with probability $\pi_{i\in j}$. The sample size is held constant at $1,000$. As the outcome is binary, we choose to fit a binomial likelihood of counts of $y_{i\in j}=1$ per cell $j$. For simplicity we fit a subset of the models used by Kuh et al. (2023), focusing on the following models:
  • Full model: $\mbox{Pr}(y_{i\in j}=1|n_{j})=\operatorname{logit}^{-1}(\beta_{0}+\alpha^{(X_{1})}_{j}+\alpha^{(X_{2})}_{j}+\alpha^{(X_{3})}_{j}+\alpha^{(X_{4})}_{j}),$
  • Precision variable model: $\mbox{Pr}(y_{i\in j}=1|n_{j})=\operatorname{logit}^{-1}(\beta_{0}+\alpha^{(X_{1})}_{j}+\alpha^{(X_{2})}_{j}+\alpha^{(X_{3})}_{j}),$
  • Bias variable model: $\mbox{Pr}(y_{i\in j}=1|n_{j})=\operatorname{logit}^{-1}(\beta_{0}+\alpha^{(X_{1})}_{j}+\alpha^{(X_{3})}_{j}+\alpha^{(X_{4})}_{j}),$
  • Nuisance variable model: $\mbox{Pr}(y_{i\in j}=1|n_{j})=\operatorname{logit}^{-1}(\beta_{0}+\alpha^{(X_{1})}_{j}+\alpha^{(X_{3})}_{j}),$
where $\alpha^{(k)}_{j}\sim\operatorname{normal}(0,\sigma^{(k)}$), $\sigma^{(k)}\sim t(3,0,2.5)$ and $\beta_{0}\sim t(3,0,2.5)$, which are the standard priors in the brms R package (Bürkner, 2017, 2018).
@@
    Create a sample from this population so that each cell has at least one individual observed, and then according to $\pi_{i\in j}$.
@@
      An estimate for each cell and compare to the true $\mbox{Pr}(Y_{i\in j}=1)$ to calculate the cellwise CRPS and squared error as per (4) and (5).
The cellwise score using the population cellwise truth is equivalent, up to a small amount of noise, to the population score; see supplementary materials for visualisation of results). However, in practice the population truth (cellwise or at the aggregate level) is not available.
One option in this case is to approximate the population truth with the sample observation, $\mbox{Pr}(y_{i\in j}|n_{j},y_{i\in j})$. This leads to
$\widehat{\operatorname{Error}}_{\mathrm{MRP}}=\frac{\sum_{j=1}^{J}N_{j}(\mbox{Pr}(Y_{i\in j}|y,\hat{\theta})-\mbox{Pr}(y_{i\in j}|n_{j},y_{j}))}{\sum_{j=1}^{J}N_{j}},$ (9)
and
$\widehat{\operatorname{CRPS}}_{\mathrm{MRP}}(y,n,\hat{\theta},\mathbf{x})=\frac{1}{2}\frac{1}{B}\sum_{b}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}\Big{(}\mbox{Pr}(Y_{i\in j}|y,\hat{\theta}^{b})-\mbox{Pr}(Y_{i\in j}|y,\hat{\theta}^{\prime b})\Big{)}}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}-\\ \frac{1}{B}\sum_{b}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}\Big{(}\mbox{Pr}(Y_{i\in j}|y,\hat{\theta}^{b})-\mbox{Pr}(y_{i\in j}=1|n_{j},y_{j})\Big{)}}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}.$ (10)
When there is only one observation per cell, the cellwise probability estimate is either 00 or $1$ accordingly. Previous research suggests that this will underestimate the magnitude of the score, which we also find (see supplementary materials for a visual demonstration). However, the ordering of models is preserved—namely, the separation between models that contain the bias variable as a predictor and models that contain only the precision variable as a predictor. The standard sum of individual scores does not maintain this ordering (e.g., see Kuh et al., 2023 for sum of ELPD, and supplementary materials for mean of CRPS and squared error scores), which suggests the scoring rule could be useful for ordering models but not estimating the expected magnitude of error.
@@ 3.2 Brute-force cross validation
To resolve the underestimation of the magnitude of scores we consider a cross validation approach. There is a wide and established literature on cross validation for estimating predictions. We considered three potential schemes: $K$-fold cross validation, leave-one-cell-out (LOCO) and leave-one-level-out (LOLO).
@@
$K$-fold cross validation partitions the sample into $K$ non-overlapping folds. The $k^{th}$ fold is removed to fit the model and score calculated using this fold. To use this approach in with MRP, we would need to adjust the score estimated in this fold to population. To do this we would need at least $K$ observations present in every poststratification matrix cell (one for each fold) which isn’t plausible given the size and complexity of survey data and the set of potential adjustment variables.
Leave-one-cell-out cross validation uses a similar idea, but instead partitions the data so each cell is it’s own fold. The $j^{th}$ cell is removed, the model fit and the score estimated using the observations in the removed cell $\mbox{Pr}(y_{j}=1|n_{j})$. In this scenario, only one observation is required per cell (to estimate the probability), and the estimate of population error can be made with our proposed score decomposition. This is the method that we will focus on in this manusucript.
Despite this, leave-one-level-out (LOLO) cross validation could be useful specifically in subpopulation estimation (see Section 4). In this instance the data are partitioned by the level of a adjustment variable and a similar process to the other validation techniques applied. In general though, it is not clear which variable to focus on when multiple variables are in the model when assessing a whole population estimate.
To use LOCO, we propose replacing our estimate $\mbox{Pr}(Y_{i\in j}=1|y,\hat{\theta})$ with $\mbox{Pr}(Y_{i\in j}=1|y_{i\notin j},\hat{\theta})$, the probability of $y=1$ in that cell given that it was not used when the model was fit to get
@@
$\operatorname{Error}^{(\mathrm{LOCO-CV})}=\frac{\sum_{j=1}^{J}N_{j}(\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta})-\mbox{Pr}(y_{i\in j}|n_{j},y_{j}))}{\sum_{j=1}^{J}N_{j}},$ (11)
and
@@
$\operatorname{CRPS}^{(\mathrm{LOCO-CV})}(y,n,\hat{\theta})=\frac{1}{2}\frac{1}{B}\sum_{b}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}\Big{(}\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}^{b})-\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}^{\prime b})\Big{)}}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}-\\ \frac{1}{B}\sum_{b}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}\Big{(}\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}^{b})-\mbox{Pr}(y_{i\in j}|n_{j},y_{j})\Big{)}}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}.$ (12)
From here on in this manuscript we will only be discussing an aggregate MRP score and so no longer note this for notation readability.
@@
When the model is refit excluding each cell (leading to $J+1$ model fits), we call this a brute-force implementation of leave-one-cell-out (LOCO). It is not efficient in terms of time and computational resources, but it does allow us to directly investigate whether the underestimation impact of cross validation on the magnitude of score estimate. Figure 1 shows that the score estimated with a brute-force LOCO approach is indeed slightly larger than that estimated using the full sample to model and estimate the score. However, this is still a considerable underestimate for the true score, as shown by Figure 2.
@@
Figure 1: Comparison of the brute-force leave-one-cell-out approximation (LOCO, $x$-axis) with a non cross validation scheme ($y$-axis, top row) and approximate cross validation scheme ($y$-axis, bottom row). Colour of point represents model, points within this represent different simulation iterations. The black line represents an identity line. We plot (-)CRPS(left panels) and squared error (right panels).
Figure 2: Comparison of a direct score of the MRP estimate (using known population truth, $y$-axis) against an indirect score of the MRP estimate (using a brute-force leave-one-cell-out approach, $x$-axis). Colour of point represents the different models fitted, points within this different simulation iterations. The black line represents an identity line. We plot squared error and the negative of CRPS, so higher can be interpreted as a worse model. These results demonstrate that using LOCO still underestimates the model score.
@@
Despite the underestimate of the magnitude of the score, we believe that this approach is still useful to practitioners who wish to evaluate the relative efficacy of different models when estimating the population mean. Unlike previous attempts, ordering is maintained. We advise that the magnitude of the score be interpreted with caution.
@@ 3.3 Approximate cross validation
The magnitude of the score is not the only challenge for a brute-force leave-one-cell-out approach. A practical issue is the sheer amount of computational required for a brute-force approach. For example, in the first iteration of the previously described simulation study, the posterior inference had to be rerun $294$ times (once for each cell plus one for full data), and this example is a relatively small toy problem in terms of the size of the poststratification matrix. Many MRP examples involve poststratifying to an area level variable like postcode, which might have $100$ levels (Machalek et al., 2022). In this section, we apply an approximate LOO method based on Pareto smoothed importance sampling (PSIS; Vehtari et al., 2017, 2024).
PSIS-LOO (equation (10) of Vehtari et al., 2017) estimates the sum expected log individual predictive density by
@@
where $w_{i}^{b}$ is the Pareto smoothed importance weight for the $b^{th}$ posterior draw (denoted $s$ by Vehtari et al., 2017) and the $i^{th}$ individual. We modify this for our purposes by using other scoring metrics. We begin with substituting the error in each cell as before, with for shorthand we denote $\epsilon_{j}^{b}=\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}^{b})-\mbox{Pr}(Y_{i\in j}|n_{j},y_{j})$. For each cell $j$ we can compute Pareto smoothed weights $w_{j}^{b}$ for each posterior sample $b$ to estimate $\epsilon_{j}^{b}$ as outlined by Vehtari et al. (2017). Using these, we can approximate $\epsilon_{j}^{b}$ and substitute into (11) as
$\widehat{\operatorname{Error}}^{(\mathrm{LOCO-CV})}=\frac{1}{N}\sum_{j=1}^{J}N_{j}\frac{\sum_{b=1}^{B}w_{j}^{b}\epsilon_{j}^{b}}{\sum_{b=1}^{B}w_{j}^{b}}$ (13)
@@
It is slightly more complicated to achieve this with the CRPS, but we begin by denoting the $g(y_{i\notin j},\hat{\theta}^{b},\hat{\theta}^{\prime b})=\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}^{b})-\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}^{\prime b})$ and $h(y_{i\notin j},y_{j},n_{j},\hat{\theta}^{b})=\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}^{b})-\mbox{Pr}(y_{i\in j}|n_{j},y_{j})$. This means we can write (12) more concisely as
$\begin{split}\operatorname{CRPS}^{(\mathrm{LOCO-CV})}(y,n,\hat{\theta},\hat{\theta}^{\prime})=\frac{1}{2}\frac{1}{B}\sum_{b=1}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}g(y_{i\notin j},\hat{\theta}^{b},\hat{\theta}^{\prime b})}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}-\\ \frac{1}{B}\sum_{b=1}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}h(y_{i\notin j},y_{j},n_{j},\hat{\theta}^{b})}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}.\end{split}$
While it is possible to simply take an importance-weighted sum for the squared error, CRPS doesn’t have an amenable form. To apply the PSIS weights, which are the $b^{th}$ weight for the $j^{th}$ cell, we do importance resampling of the posterior draws for each cell using a stratified resampling algorithm (Kitagawa, 1996) using the $B$ PSIS weights for each cell. This procedure is implemented in the posterior R package (Bürkner et al., 2023). This produces $B$ resampled posterior draws $\hat{\theta}_{RW}$, which we permute to calculate PSIS reweighted $\hat{\theta}_{RW}^{\prime}$, and thus $g_{j}^{\prime}(\hat{\theta}_{RW}^{b},\hat{\theta}_{RW}^{\prime b})=\mbox{Pr}(Y_{i\in j}|\hat{\theta}_{RW}^{b})-\mbox{Pr}(Y_{i\in j}|\hat{\theta}_{RW}^{\prime b})$ and $h_{j}^{\prime}(y_{j},n_{j},\hat{\theta}_{RW}^{b})=\mbox{Pr}(Y_{i\in j}|\hat{\theta}_{RW}^{b})-\mbox{Pr}(Y_{i\in j}|n_{j},y_{j})$, to get
$\widehat{\operatorname{CRPS}}^{(\mathrm{LOCO-CV})}(y,n,\hat{\theta}_{RW},\hat{\theta}_{RW}^{\prime})=\\ \frac{1}{BN}\Bigg{(}\frac{1}{2}\sum_{b=1}^{B}\Bigg{(}\Bigg{\lvert}\sum_{j=1}^{J}N_{j}g_{j}(\hat{\theta}_{RW}^{b},\hat{\theta}_{RW}^{\prime b})\Bigg{\rvert}-\Bigg{\lvert}\sum_{j=1}^{J}h_{j}(y_{j},n_{j},\hat{\theta}_{RW}^{b})\Bigg{\rvert}\Bigg{)}\Bigg{)}.$ (14)
The second row of Figure 1 shows that the approximate PSIS-LOCO scores are approximately equivalent to the brute-force LOCO scores.
@@
When we followed the traditional approach of comparing the mean of cell scores to the true mean of cell scores, the leave-one-cell-out approach appeared to slightly over estimate the magnitude of score (see supplementary), suggesting that the observed underestimation is a feature of our particular score decomposition.